OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 96π in3 27π in3 65π in3 105π in3

OpenStudy (btaylor):

\[V(vase)=(12)(3^2)\pi=108\pi\]\[V(marbles)=9(\frac{4\pi(1^3)}{3}) = 12 \pi\]\[V(water)=V(vase)-V(marbles) = 108 \pi - 12 \pi = 96 \pi\]

OpenStudy (anonymous):

I looked at another example and i couldnt find the answer! Thanks for the help you deffinately helped a lot!

OpenStudy (btaylor):

yw